• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Deur

    Log entry time 21:59:29 on October 6,2002

    Entry number 87500

    keyword=Setup for kinematics ep4-4

    HRS left is now at 1.687 GeV/c
    HRS right is now at 3.064 GeV (Q2 and Q3 cycled)
    We moved left HRS to 43.604
    We moved right HRS to 22.318
    We opened the 1.5" and 3" FPP carbon doors